# The Loser’s Curse Revisited: A Case Study of the Critical Role of Specifying a Utility Function

### v2

@@ Appendix C Estimating the trade market value of draft picks
The aforementioned assumptions range from quite strong to very strong and deserve to be examined. The fourth assumption, that ${v_{\text{market}}}$ follows a Weibull distribution, seems to be the most mild, as the Weibull is a sufficiently expressive family that can capture many different plausible draft value curves. The first assumption, additivity, is wrong but mild. For instance, the value of a bundle of $12$ picks is not additive because not all of these players can play on the field at the same time. But additivity doesn’t seem to be a terrible assumption because most trades involve just a few picks. The third assumption is wrong; they probably meant to assume that observed trades are fair on average:
@@
Despite these issues with the simplifying assumptions, we replicate the analysis discussed above from a dataset consisting of all draft trades consisting purely of draft picks from 2006-2023. In light of the aforementioned flaws, we use the fitted ${v_{\text{market}}}$ curve to tell a general story about the steepness of the trade market. We take it with a grain of salt without obssessing over the exact shape of the curve or its minutae. We display the ${v_{\text{market}}}$ from Massey and Thaler, (2013) and from our replication, with and without the discount factor, in Figure 18.
Figure 18: Weibull trade market value curves ${v_{\text{market}}}$ according to Massey and Thaler, (2013) and our replication.

### v3

@@ Appendix C Estimating the trade market value of draft picks
The aforementioned assumptions range from quite strong to very strong and deserve to be examined. The fourth assumption, that ${v_{\text{market}}}$ follows a Weibull distribution, seems to be the most mild, as the Weibull is a sufficiently expressive family that can capture many different plausible draft position value curves. The first assumption, additivity, is wrong but mild. For instance, the value of a bundle of $12$ picks is not additive because not all of these players can play on the field at the same time. But additivity doesn’t seem to be a terrible assumption because most trades involve just a few picks. The third assumption is wrong; they probably meant to assume that observed trades are fair on average:
@@
Despite these issues with the simplifying assumptions, we replicate the analysis discussed above from a dataset consisting of all draft trades consisting purely of draft picks from 2006-2023. In light of the aforementioned flaws, we use the fitted ${v_{\text{market}}}$ curve to tell a general story about the steepness of the trade market. We take it with a grain of salt without obssessing over the exact shape of the curve or its minutae. We display the ${v_{\text{market}}}$ from Massey and Thaler, (2013) and from our replication, with and without the discount factor, in Figure 18. These curves share the same general shape––a monotonically decreasing and extremely steep trend––indicating general managers’ stark preference for top picks. Value declines rapidly, falling to between one-half and one-quarter of a first pick by the middle of the first round. Value then decays to less than a tenth of a first pick by the start of the second round before quickly plummeting to near-zero thereafter.
Figure 18: Weibull trade market value curves ${v_{\text{market}}}$ according to Massey and Thaler, (2013) and our replication. Massey and Thaler’s curves here do not use a discount factor.